OpenStudy (anonymous):

Solve the following equation: 8m^2 = 14 − 9m

OpenStudy (anonymous):

rewrite the equation first

OpenStudy (anonymous):

8m^2+9m-14=0?

OpenStudy (anonymous):

yeah, now your factors are gonna look a lil different, they will have coefficients on them this time (4x )(2x ) or (8x )(x )

OpenStudy (anonymous):

it'll be 7 & 2, but I don't know which is negative.

OpenStudy (anonymous):

I choose 4x & 2x

OpenStudy (anonymous):

okay plug 7 and 2 and see if it works.

OpenStudy (anonymous):

you're not wrong, 7 and 2 are the constants you'll be using but the 4x & 42 won't work, you have to use the 8x and x

OpenStudy (anonymous):

-7 & 2

OpenStudy (anonymous):

this kind of problem is basically try all possible ways until u find the correct way.. ill save you the trouble the factors are (8m-7)(m+2)=0

OpenStudy (anonymous):

That's what I have so far, now I have to subtract & stuff.

OpenStudy (anonymous):

to find m now you have to set each factor equal to 0. 8m-7=0 m+2=0

OpenStudy (anonymous):

and solve for

OpenStudy (anonymous):

m

OpenStudy (anonymous):

7/8 & -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you, next quesion.

OpenStudy (anonymous):

lmao are you doing semester 1 of algebra 2 on flvs ?

OpenStudy (anonymous):

I don't think so lol